OpenStudy (anonymous):

Solve the following quadratic equation using the quadratic formula and then choose the correct solution set. 5x2 + 7x = 6 {-2, 2} {-2, } {-, 2}

OpenStudy (anonymous):

(-2.0.6)

OpenStudy (anonymous):

This is the formula they use. Can you use it and work it step by step?

OpenStudy (anonymous):

\[x _{1,2}=-7\pm \sqrt{7^{2}-4*5*(-6)}/2*5=-7 \pm \sqrt{49+120}/10=(-7\pm 13)/10=(-2,3/5)\]

OpenStudy (anonymous):

{-2,3/5}